OpenStudy (anonymous):

integral from 0 to pi/2 of sin^5 x cos^12 x dx without using the reduction formula

OpenStudy (anonymous):

@SithsAndGiggles you did good in the last one, i did get it solved, not by the way you did it with half-angle formula but i used your input to check my answer. thank you.

OpenStudy (anonymous):

\[\int_0^{\pi/2}\sin^5x\cos^{12}x\;dx\] I would start by \[\begin{align*}\sin^5x&= \sin x \sin^4x\\ &=\sin x (\sin^2x)^2\\ &=\sin x (1-\cos^2x)^2\end{align*}\] Expand that, and you get something like \[\int_0^{\pi/2}\left(\sin x \cos^{12}x - 2\sin x \cos^{14}x + \sin x\cos^{16}x\right)\;dx\] (I didn't double-check that, so it may be wrong) Then a simple u-sub shouldn't be too difficult.

OpenStudy (kainui):

Technically you can do this in the same way that the reduction formula is derived by going through a very long and tedious integration by parts. I'm trying to figure out if there's some clever way to cancel it out with one sweeping u-substitution or something since the derivative of sine is just cosine. Yeah that looks pretty alright to me.

OpenStudy (anonymous):

yeah, @Kainui, that is what i would like to use, but my prof. wants us to use the very long and tedious integration by parts... that is how it will be on the exam in a week or so

OpenStudy (anonymous):

@SithsAndGiggles, good that is were I started. I'm going to keep working and see what happens

OpenStudy (anonymous):

Did you say you wanted to do this by parts? I don't think there's any good way to do that.

OpenStudy (anonymous):

Actually, what I meant to say was U-sub.

OpenStudy (anonymous):

Just to make sure does \[(1-\cos^2 x)^2 = (1-2\cos^4 x+\cos^4 x) \]?

OpenStudy (anonymous):

Close, the power of the second cosine term should be 2.

OpenStudy (anonymous):

\[(1−2\cos^2x+\cos^4x)\] ? Now also how did you expand from here? I know you plug in the \[sinx(1−\cos^2x)^2 \] for \[\sin^5x\] but how to expand from there?

OpenStudy (anonymous):

oh haha nevermind. sorry.

OpenStudy (anonymous):

well, @SithsAndGiggles what do you suggest i use for u-sub, just using sinx and break it down multiple times?

OpenStudy (anonymous):

Let u = cosx, -du = sinx dx. Then you have \[\left(-\int_1^0u^{12}du\right) + \left(-\int_1^0u^{14}du\right) + \left(-\int_1^0u^{16}du\right)\\ \int_0^1u^{12}du+\int_0^1u^{14}du+\int_0^1u^{16}du\]

OpenStudy (anonymous):

wouldn't the \[\int\limits u^{12}\] be negative?

OpenStudy (anonymous):

Well here is the other thing too... This is a definite integral too... From 0 to pi/2 but I'm trying to solve the indefinite one first then deal with plugging in the numbers after

OpenStudy (anonymous):

Sorry, I'm juggling a few things at a time here. Are you confused about the limits on the integrals? By changing integration over [1,0] to [0,1], the sign changes.

OpenStudy (anonymous):

\[\int\limits -u^{12}+2u^{14}-u^{16}du\]

OpenStudy (anonymous):

indeed, but if i dont change the limits then i can set it like this?

OpenStudy (anonymous):

\[[-\frac{ u^{13} }{ 13 }+\frac{ 2u^{15} }{ 15 }-\frac{ u^{17} }{ 17 }]\]

OpenStudy (anonymous):

Uh, let me rewrite what I've done. I feel like I'm tripping myself up a bit here... \[\int_0^{\pi/2}\left(\sin x\cos^{12}x - \sin x\cos^{14}x + \sin x\cos^{16}x\right)dx\\ \int_0^{\pi/2}\sin x\cos^{12}x\;dx - \int_0^{\pi/2}\sin x\cos^{14}x\;dx + \int_0^{\pi/2}\sin x\cos^{16}x\;dx\] After the sub, we get \[\left(-\int_1^0u^{12}\;du\right) - \left(-\int_1^0u^{14}\;du\right) + \left(-\int_1^0u^{16}\;du\right)\\ \int_0^1u^{12}\;du-\int_0^1u^{14}\;du + \int_0^1u^{16}\;du\] So it looks like I had a single sign mistake. Do you follow this?

OpenStudy (anonymous):

Yeah, that is where I'm at right now. I am subbing in cosx for u right now and then solving from 0 to pi/2.

OpenStudy (anonymous):

oh my - i came out to 0... that doesnt seem right

OpenStudy (anonymous):

Well, before you do that (over 0 to pi/2, I mean), make sure you sub back u = cosx. Otherwise you'll get a wrong answer. I also left out a constant 2 on the second integral... Sorry about that.

OpenStudy (anonymous):

Sooo this is what i just had written down... \[[-\frac{ \cos^{13}x }{ 13 }+\frac{ 2\cos^{15}x }{ 15 }-\frac{ \cos^{17}x }{ 17 }]\int\limits_{0}^{\pi/2}\] now that is not integral that is just meaning to plug in the endpoints into the equation.

OpenStudy (anonymous):

When I just plugged this into my calculator, I came out with 0 though... I also tried plugging it into my homework and it did not accept it.

OpenStudy (anonymous):

I guess the answer should be 0.002413273 when using the solver on my calculator.

OpenStudy (anonymous):

Via the FTC, that last expression you have becomes \[-\frac{1}{13}\left(\cos^{13}\frac{\pi}{2} - \cos^{13}0\right)+\frac{2}{15}\left(\cos^{15}\frac{\pi}{2} - \cos^{15}0\right)-\frac{1}{17}\left(\cos^{17}\frac{\pi}{2} - \cos^{17}0\right)\] Which is equal to \[-\frac{1}{13}(0-1)+\frac{2}{15}(0-1)-\frac{1}{17}(0-1)\\ =\frac{1}{13}-\frac{2}{15}+\frac{1}{17}\\ =\frac{8}{3315}\\ \approx0.00241\]

OpenStudy (anonymous):

I feel so silly, when I plugged it into the calc I only did pi/2 i forgot that cos0 =1 so it wouldn't end up being all zero'd out for the second part... Thank you for walking through this with me.

OpenStudy (anonymous):

I don't know if I was thinking sinx or what. either way - very good job @SithsAndGiggles !!!

OpenStudy (anonymous):

Thanks :3